OpenStudy (meilendurcer):

Determine which ordered pair is a solution of y = x^2– 2. A. (–2, 6) B. (5, 21) C. (–1, –1) D. (–5, 24)

OpenStudy (anonymous):

Plug in each ordered pair into the equation and see which ones hold true. For example with an ordered pair like \((1, 0)\), we'd plug it into \(y = x^2-2\) like so: \[ (1) = (0)^2 -2 \\ 1 = 0-2 \\ 1 = -1 \]Since it's obvious that \(1 \neq -1\), the ordered pari \((1, 0)\) is not a solution.

OpenStudy (anonymous):

The answer is C (-1,-1).

OpenStudy (meilendurcer):

thanks.

OpenStudy (anonymous):

welcome

OpenStudy (meilendurcer):

can you help with another?